Your firm has limited capital to invest and is therefore interested in comparing projects based on the profitability index (PI), as well as other measures. What is the PI of the project with the estimated cash flows below? The required rate of return is 18.2%. Round to 3 decimals.

Year 0 cash flow = -710,000
Year 1 cash flow = -190,000
Year 2 cash flow = 370,000
Year 3 cash flow = 490,000
Year 4 cash flow = 510,000
Year 5 cash flow = 460,000
Answer #1

D9 fx -SUM(D3:08) 2 year Cash flows 3 0 $ 1 $ 5 2 $ pv@18.2% Present value (7,10,000.00) 1.0000 $ (7,10,000.00) (1,90,000.00)

D9 for =SUM(D3:08) А 1 2 year 30 pv@18.2% Cash flows -710000 -190000 370000 490000 510000 460000 =C3/1.182 =C4/1.182 =C5/1.18